OpenStudy (amonoconnor):

How would one go about taking the derivative of: "2π" (?) Any and all help is greatly appreciated! :)

OpenStudy (cggurumanjunath):

0

OpenStudy (cggurumanjunath):

2 and pi are constants !

OpenStudy (cggurumanjunath):

derivative of a constant is zero !

OpenStudy (anonymous):

derevative of any constant is 0

OpenStudy (isaiah.feynman):

One of the easiest questions Calculus can ask.

zepdrix (zepdrix):

Try not to get confused by these constants dressed up in their fancy tuxedos. They're just trying to trick you.\[\Large \frac{d}{dx}e^3\quad=\quad 0\]\[\Large \frac{d}{dx}\ln(2)\quad=\quad 0\]\[\Large \frac{d}{dx}2\pi\quad=\quad 0\]There is no variable term present in any of these, so they're all constant.

OpenStudy (amonoconnor):

Thanks for the explanation @zepdrix. You were spot on- I was getting thrown off by the π. Thanks again !

OpenStudy (unklerhaukus):

\[\frac{\mathrm d}{\mathrm dx}(2\pi)=\frac{\mathrm d}{\mathrm dx}(2\pi\cdot x^0)=0\times 2\pi \cdot x^{-1}=0\]

OpenStudy (amonoconnor):

Thank you! :)